UT 20150120 LUCI2 AO Daytime Commissioning

Observer LUCI-Team: WSeifert
Observer AO-Team: SEsposito, GAgapito, EPinna, APuglisi
Telescope Operator:
LBTO Support:

Plan for the day

- Installation of ND-mount

Summary

- Installation of ND-mount successfull

- MOS-issue (robot grabber rotation) but solved

Details

MOS:

there has been in the night before an error for the motion of M5 (robot grabber rotation)
-> Peter and James R made measurements and found the amplifier current too low;
by increasing to nominal value the MOS worked again.

ND filter bracket installation

LUCI2 dismounted 8:30
Filter mount and bracket installed
instrument re-installed 12:30
pumped until 14:45, no improvement by the pumping
cooling started 15:00

LUCI2_AGWFilter_20150121.jpg

Image of AO ND-filter in AGw taken with N3.75-camera:

No AO related measurements today

NCP correction discussion

request for bad pixel map
request for linearity coeffs
request for subwindow read-out

All times are UT
Files are luci2.20150120.NNNN.fits

-- %USERSIG{JochenHeidt - 2015-01-21}%
I Attachment Action Size Date Who Comment
LUCI2_AGWFilter_20150121.jpgjpg LUCI2_AGWFilter_20150121.jpg manage 84 K 22 Jan 2015 - 09:37 JochenHeidt Image of AO ND-filter in AGw taken with N3.75-camera
Topic revision: r5 - 30 Jan 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback